OpenStudy (anonymous):

An urn contains six red balls, six white balls, and six blue balls, and sample of four balls is drawn at random without replacement. Compute the probability that all of the balls in the sample are the same color. (Round your answer to four decimal places.)

OpenStudy (welshfella):

basically you need to calculate P(6 red or 6 blue or 6 white)

OpenStudy (anonymous):

yea pretty much in simpleest form it is what @welshfella said im not gonna go further in depth and get you confused.

OpenStudy (anonymous):

Could you please bare with me and go further in steps to help me solve this. I was totally lost thats why I am here seeking help. Thank you.

OpenStudy (welshfella):

to find probability of drawing 4 reds you multiply the probabilities P(red on first draw) = 6/18 = 1/3 then P(red on 2nd draw) = 5/17 do you follow that ok?

OpenStudy (welshfella):

- then you find probablity of red on third and fourth draw in a similar way

OpenStudy (welshfella):

now because these draws are independent to find probability of getting 4 reds you multiply them

OpenStudy (welshfella):

so P( 4 reds) = 1/3 * 5/17 * A * B) where A and B are what you get for draws 3 and 4

OpenStudy (welshfella):

have you any questions on the above posts?

OpenStudy (anonymous):

lol well it seems @welshfella has this problem . Hope you understand ;) @amool

OpenStudy (anonymous):

Thank you so much. I understand much better now.

OpenStudy (welshfella):

good

OpenStudy (anonymous):

ok sorry how did you get the 5/17.

OpenStudy (welshfella):

originally there are 18 balls so after 1 red is taken out there are 5 reds left and a total of 17 balls ( as the balls are not replaced)

OpenStudy (welshfella):

So for the third draw there are only 4 reds and a total of 16 balls. so P(red on 3rd draw) = 4/16 = 1/4

OpenStudy (welshfella):

so for 4th draw what is P(red ball)?

OpenStudy (anonymous):

so it will be 3/15 = 1/5 I think this will be the 4th draw

OpenStudy (welshfella):

right

OpenStudy (anonymous):

Thanks for baring with me now I fully know the method thank you.

OpenStudy (welshfella):

When you have AND involved with probabilities think Multiply red ANd red ANd red AND red so P( 4 reds) = 1/3 * 5/17 * 1/4 * 1/5 which you can convert to a decimal as they want decimals in the answer

OpenStudy (anonymous):

yeah it is .004902

OpenStudy (welshfella):

yes Now the probabilities of getting 4 blues and 4 whites is exactly the same both 0.004902

OpenStudy (welshfella):

Noe here we have an OR involved red OR Blue OR white In this case you ADD the probabilities (OR = ADD) so final answer is the sum of the 3 probabilities

OpenStudy (welshfella):

so ts 3 * 0.004902

OpenStudy (anonymous):

got it .014706